OpenStudy (anonymous):

HELP!!! Simplify the fraction: 10 over the square root of 2

OpenStudy (jdoe0001):

divide top and bottom by 2

OpenStudy (anonymous):

multiple choice: A. five-halves times the square root of 2 B. 5 times the square root of 2 C. 10 times the square root of 2 D. 5 is it B?? since u reduce all it down??

OpenStudy (anonymous):

@jdoe0001 ???

OpenStudy (darkigloo):

yes it is b

OpenStudy (jdoe0001):

well.. what did you get from the division?

OpenStudy (jdoe0001):

\(\bf \cfrac{10\div 2}{2\div 2}\)

OpenStudy (anonymous):

so its 5 @jdoe0001

OpenStudy (darkigloo):

its not 5.. \[\frac{ 10 }{ \sqrt{2} } \] multiply the top and bottom by sqrt 2 \[\frac{ 10\sqrt{2} }{ \sqrt{2}*\sqrt{2} }\] \[\frac{ 10\sqrt{2} }{ 2 }\]

OpenStudy (darkigloo):

divide by 2 on top and bottom

OpenStudy (darkigloo):

5√2 is your answer.

OpenStudy (anonymous):

now im torn between u and jdoe lol

OpenStudy (darkigloo):

i'm not sure why jdoe said divide by 2.

OpenStudy (darkigloo):

**Whenever you have a radical in the bottom of a fraction, you multiply the top and bottom by that radical number.